How does multiplying both dividend and the divisor by factor 10 sometidas makes a promblem easier to solve
1
Expert's answer
2019-11-24T11:59:48-0500

The reason: if we deal with decimals with 1 number after the decimal point, multiplying both numbers will make these both numbers integer and then these integers could be simplified in an easier way; this will not affect the final answer.

Example:

1.5 : 2.5=?

Multiplying both numbers by 10

2.5 * 10 = 25

1.5 * 10 = 15

Then

1.5 : 2.5 = 15 : 25 = (5 * 3) : (5 * 5) = 3 : 5 = 0.6